# Human Perception-Guided Pretraining for Increased Generalization

### v2

@@ Abstract
Incorporating human perception into training of convolutional neural networks (CNN) has boosted generalization capabilities of such models in open-set recognition tasks. One of the active research questions is where (in the model architecture) and how to efficiently incorporate always-limited human perceptual data into training strategies of models. In this paper, we introduce MENTOR (huMan pErceptioN-guided preTraining fOr increased geneRalization), which addresses this question through two unique rounds of training the CNNs tasked with open-set anomaly detection. First, we train an autoencoder to learn human saliency maps given an input image, without class labels. The autoencoder is thus tasked with discovering domain-specific salient features which mimic human perception. Second, we remove the decoder part, add a classification layer on top of the encoder, and fine-tune this new model conventionally. We show that MENTOR’s benefits are twofold: (a) significant accuracy boost in anomaly detection tasks (in this paper demonstrated for detection of unknown iris presentation attacks, synthetically-generated faces, and anomalies in chest X-ray images), compared to models utilizing conventional transfer learning (e.g., sourcing the weights from ImageNet-pretrained models) as well as to models trained with the state-of-the-art approach incorporating human perception guidance into loss functions, and (b) an increase in the efficiency of model training, requiring fewer epochs to converge compared to state-of-the-art training methods.
@@ 1 Introduction
Figure 1: MENTOR approach. Step 1: An autoencoder-based model is first trained to recreate human saliency maps, and thus to build an understanding of human perception-sourced salient features into the encoder $\mathcal{F}_{\text{encoder}}$. Step 2: Such pre-trained encoder $\mathcal{F}_{\text{encoder}}$ is then decoupled from the autoencoder, and along with a classifier $\mathcal{F}_{\text{class}}$ are tuned for the anomaly detection task utilizing standard cross-entropy loss.
@@
Human perceptual information is often incorporated into deep learning training strategies in order to improve generalization [3, 33], align models with human-sourced saliency [5, 10], and reduce training time by supplying additional prior knowledge into the training process [5, 43]. However, a common and valid criticism is the high cost of acquiring human annotations or eye tracking-sourced saliency data. For anomaly detection tasks, such as biometric presentation attack detection (PAD), new attack types are developed frequently. Medical diagnostics may require continual update of human perceptual understanding of anomalies present in medical samples. In all these tasks, requiring few-shot learning solutions functioning well in an open-set classification regime, collecting more training samples and/or more human saliency data is difficult or even impossible. This is why an effective use of the existing, although always-limited human saliency information originating from domain experts is crucial. An associated question is how and where to appropriately incorporate human perceptual understanding of a task into model training to deliver strong priors allowing for better generalization of such models.
@@
This paper introduces MENTOR, a framework for human perception-guided pretraining, evaluated in the context of open-set iris PAD, synthetic face detection, and disease classification from chest X-ray scans. MENTOR relies on the intuition that (as children) we first learn the representations of the visual world without knowing the “class labels,” and we assign meaning in what we see later in our lives. Around being one year old, infants began to point, imitate, and understand objects merely by observation or visual salience, without any explicit knowledge of the object’s application or type (classification) [16]. We emulate this iterative process by first training an autoencoder to learn the associations between an input image and a human saliency map (only visual salience, with no class labels). Once this representation is learned, the decoder-side is removed, a classifier is put on top of the encoder’s latent space, and the entire model is then fine-tuned in a regular way using cross-entropy loss to solve a classification task at hand. Fig. 1 illustrates this pipeline. Our experiments show that such class label-agnostic, but human perception-aware pretraining of the model’s backbone allows for (a) better generalization to unknown samples than models fine-tuned in the same way but initialized with weights obtained in pre-training on a massive amount data such as ImageNet, (b) same as (a) when compared to models trained with a state-of-the-art human perception guidance incorporated into a loss function, and (c) faster convergence compared to models initialized with ImageNet weights. We make these observations for three different domains, in which humans can deliver meaningful saliency information (iris PAD, synthetic face detection, and chest X-ray-based diagnosis), and for three different neural network architectures (ResNet, Inception and EfficientNet). In other words, MENTOR efficiently incorporates the limited human perceptual understanding of a domain into training of convolutional neural networks (CNN).
@@
Two core differences between MENTOR and state-of-the-art unsupervised representation learning approaches, such as DINO [6], are that (a) MENTOR leverages human perception-based guidance in the pre-training phase, and (b) MENTOR operates in a very data-limited regime, since it’s targeted to domains, in which collecting data is intricate, but humans competent in a given domain can be found. It is also noteworthy that MENTOR does not assume any particular CNN architecture, and works without any architectural changes. For clarity, this paper is organized around the following two research questions:
@@
  • RQ1: Does MENTOR improve the performance of iris presentation attack detection, synthetic face detection, and chest X-ray-based diagnosis compared to models pre-trained with massive amount of non-human-perception-sourced data (ImageNet [13])? (answered in Sec. 5.1)
  • RQ2: Does MENTOR allow for faster model’s fine-tuning, compared to pre-training with large non-human-perception-sourced data? (answered in Sec. 5.2)
The sources codes of the proposed approach are offered with this paper [^1] Codes will be released when the peer-reviewed version is published..
@@ 2.1 Human Saliency-Guided Training
The most architecture-agnostic implementations of incorporating human saliency into model’s training are those using perception-specific loss function components. Boyd et al.[5] introduced the CYBORG loss function, which penalizes the model for divergence between human saliency maps and model’s Class Activation Maps (CAM) [47]. While this method aligns the activations of the model’s last convolutional layer with human saliency, it’s unclear if that mechanism actually builds a human perceptual intelligence into the model, especially in its earlier layers, or acts as a regularizer by lowering the entropy of model’s salience [36]. The MENTOR approach proposed in this paper is different from CYBORG (and other similar human perception-based training strategies) in a sense that it encourages the model to first build associations between input images and human-sourced salient features, without telling the model what task these features are useful for, to create more general interpretations of the visual world.
@@ 2.2 Efficient Use of Human Annotations
Crum et al. [8] applied human salience in a Teacher-Student training paradigm to make a better use of limited human salience data. In their work, human annotations were first used to train “teacher” models, which are then used to generate subsequent model salience for training “student” models. Teachers were trained using the CYBORG loss, and afterwards used to generate saliency maps on un-annotated training samples using CAM [47] and RISE [34]. This training paradigm boosted performance in synthetic face detection and iris PAD. Interestingly, the MENTOR’s byproduct is an autoencoder that generates human saliency maps for unseen data, and thus it can complement approaches such as the above teacher-student learning paradigms.
@@ 3.1 Methodology
MENTOR is a novel approach of incorporating human salience into model training through a two-part training series. First, an autoencoder is trained to generate human-like saliency given an input image, but without any class labels (see “Step 1” in Fig. 1). In that way, the model creates associations between input samples and human-salient regions. Once these associations are made, we put a single, fully-connected layer (initialized with random weights) on top of the encoder’s embeddings and fine-tune the entire architecture to solve the classification task (see “Step 2” in Fig. 1).
It’s analogous to theories involving how humans gain perceptual understanding of the visual world, by first being exposed to visual stimuli without being given explanations (or “class labels”) of what they see. As shown later in Sec. 5), such human perception-based pre-training results in a better performance in three anomaly detection tasks (from three different domains) compared to those observed for models initialized with weights obtained after training with a large visual dataset (ImageNet), and than for models trained with state-of-the-art loss function-based human perception guidance.
@@
where $y_{k}$ is a class label for the $k$-th sample, $\bm{1}$ is an indicator function equal to $1$ when $y_{k}\in\mathcal{C}_{c}$, and 0 otherwise, $C$ is the total number of classes, $K$ is the number of samples in a batch, $\textbf{s}_{k,\mathcal{C}_{c}}^{\text{human}}$ and $\textbf{s}_{k,\mathcal{C}_{c}}^{\mathcal{F}}$ are the human and the model $\mathcal{F}$’s saliency for the $k$-th sample representing class $\mathcal{C}_{c}$, respectively, and finally $\mathcal{D}$ is the measure of dissimilarity between saliency maps (e.g.,the Mean Square Error distance).
@@ 3.2 Neural Network Architectural Choices
The MENTOR approach does not specify a type of autoencoder. We experimented with UNET [38] and UNET++ [48] with three different backbones: ResNet152 [17], Inception-V4 [39] and EfficientNet-b7 [40]) using the same training configurations [19] to observe stability of the proposed approach across visual domains and architectures. Specific domains and backbones for the autoencoder reported in Sec. 5 are as follows: UNET (all backbones for iris PAD domain) and UNET++ (all backbones for synthetic face detection and chest X-ray-based diagnosis).
Also, MENTOR does not specify the classifier put in Step 2 on top of the encoder embeddings. We demonstrate effectiveness of this approach by adding a single-layer (linear) classifier to minimize the extra architectural components added to human saliency-guided pre-trained encoder, but there are no theoretical reasons for not exploring the generalization gain achieved for deeper structures as a future research topic.
@@ 3.3 Training and Evaluation Details
MENTOR does not need any specific optimizer or hyperparameter settings. For reference purposes, we provide settings we have used in this work. The autoencoders in Step 1 were trained with the AdamW optimizer, with a learning rate of 0.0001, and with a batch size of 8. MSE loss was used to assess the agreement bewteen human saliency and the predicted saliency maps, all scaled to a canonical $224\times 224$ pixel resolution. Training was continued until 50 epochs, but many models converged quickly (in less than 5 epochs). All full models in Step 2 were trained using cross-entropy loss using Stochastic Gradient Descent (SGD), with a learning rate of 0.005 decreased by 0.1 every 12 epochs. As in Step 1, maximum number of epochs was 50.
We evaluate the proposed MENTOR approach using Area Under the Receiver Operating Characteristic Curve (AUROC). For all experiments we perform ten independent training runs instantiated using different seeds. We use boxplots with notches representing 95% confidence intervals of the median values for visual assessment of the observed differences between methods. For formal assessment of statistical significance of these differences, we use two-sample one-tailed Kolmogorov-Smirnov test with the null hypothesis stating the equality of mean AUROCs, and the alternative hypothesis stating better AUROC achieved by MENTOR. In addition, we also report the performance of the best three performing models (out of the ten training runs). The latter helps in estimating the highest performance that can be achieved when “cherry-picking” the models is justified (e.g.,for final deployment).
@@ 4.1 Training and Validation Data Subsets
For training the subsequent iris PAD classification models (“Step 2” in Fig. 1), the same 765 images (but now with class labels and no salience maps) were used as training images, in conjunction with additional 23,312 validation images sampled from the same superset as used in [3]. The train and validation sets are subject-disjoint.
@@
We trained the autoencoder using real and synthetically-generated face images and human saliency maps offered by [5]. 363 humans recruited via Amazon Mechanical Turk annotated regions of the image deemed important to their classification decision (real / fake). Only correctly classified samples (images and accompanying saliency maps) were used during training. For simplicity, 765 samples were randomly selected without replacement to match the number of samples available for Domain 1 (Iris Presentation Attack Detection).
@@
The autoencoder was trained with normal and abnormal chest X-ray images [21], and corresponding eye-tracking visual salience collected from radiologists examining these X-ray scans [2]. We consider only abnormal categories represented by at least 200 training images, that is: atelectasis, cardiomegaly, edema, lung opacity, pleural effusion, pneumonia, and support devices. We removed images containing erroneous class labels and consider only images where radiologists were certain of their classification decision. The intensity of pixels in eye tracking-based salience maps is proportional to the strength of the eye fixation. If a sample included human saliency maps from multiple radiologists, we combined these maps by taking the maximum intensity for each pixel. This was done to preserve salient information provided by each expert, who may consider different features when classifying the samples, instead of considering only the most “popular” features. All images were resized to $224\times 224$. For consistency, 765 samples and accompanying saliency maps were used during the training phases to match with sample sizes used in two other domains.
For training the subsequent chest X-ray classification models (“Step 2” in Fig. 1), all 765 images were used as training images, in conjunction with additional 4,878 subject-disjoint validation samples sampled from the same dataset.
@@ 5 Results
To answer both research questions, we compare MENTOR-trained models with those (a) pre-trained using a classical transfer learning approach and initialized with weights resulting from training the models with ImageNet samples, and (b) trained with a state-of-the-art human perception-guided approach (CYBORG) after initializing the weights with ImageNet-trained models’ weights. The graphical summaries (in a form of boxplots, along with numerical values for mean and standard deviations) of the accuracy for each domain-architecture combination are shown in Figs 3-5. The extra top-three-models-based results are summarized in Tab. 1. Validation accuracies as a function of epoch number for all domain-architecture combinations are shown in Figs 6-8.
Table 1: Mean and standard deviations of the top three AUROC scores (calculated independently for each combination of domain-architecture).
Dataset ImageNet ImageNet MENTOR
       Backbone + Xent + CYBORG [5] (this paper)
@@
       ResNet152 0.900$\pm$0.006 0.914$\pm$0.003 0.939$\pm$0.009
       Inception-V4 0.853$\pm$0.029 0.830$\pm$0.016 0.939$\pm$0.012
       EfficientNet-7b 0.895$\pm$0.005 0.832$\pm$0.007 0.922$\pm$0.004
@@
       ResNet152 0.579$\pm$0.007 0.568$\pm$0.045 0.601$\pm$0.011
       Inception-V4 0.766$\pm$0.016 0.743$\pm$0.005 0.789$\pm$0.006
       EfficientNet-7b 0.489$\pm$0.009 0.535$\pm$0.018 0.600$\pm$0.024

### v3

@@ Abstract
Leveraging human perception into training of convolutional neural networks (CNN) has boosted generalization capabilities of such models in open-set recognition tasks. One of the active research questions is where (in the model architecture or training pipeline) and how to efficiently incorporate always-limited human perceptual data into training strategies of models. In this paper, we introduce MENTOR (huMan pErceptioN-guided preTraining fOr increased geneRalization), which addresses this question through two unique rounds of training CNNs tasked with open-set anomaly detection. First, we train an autoencoder to learn human saliency maps given an input image, without any class labels. The autoencoder is thus tasked with discovering domain-specific salient features which mimic human perception. Second, we remove the decoder part, add a classification layer on top of the encoder, and train this new model conventionally, now using class labels. We show that MENTOR successfully raises the generalization performance across three different CNN backbones in a variety of anomaly detection tasks (demonstrated for detection of unknown iris presentation attacks, synthetically-generated faces, and anomalies in chest X-ray images) compared to traditional pretraining methods (e.g., sourcing the weights from ImageNet), and as well as state-of-the-art methods that incorporate human perception guidance into training. In addition, we demonstrate that MENTOR can be flexibly applied to existing human perception-guided methods and subsequently increasing their generalization with no architectural modifications.
@@ 1 Introduction
Figure 1: MENTOR approach. Step 1: An autoencoder-based model is first trained to recreate human saliency maps, and thus to build an understanding of human perception-sourced salient features into the encoder $\mathcal{F}_{\text{encoder}}$. Step 2: Such pre-trained encoder $\mathcal{F}_{\text{encoder}}$ is then decoupled from the autoencoder, and along with a classifier $\mathcal{F}_{\text{class}}$ are tuned for the anomaly detection task utilizing standard cross-entropy loss. MENTOR has been designed and evaluation to address generalization required in a single domain (in this paper independently for iris presentation attack, synthetic face and chest X-ray-based anomaly detection).
@@
Human perceptual information is often incorporated into deep learning training strategies in order to improve generalization [2, 34], align models with human-sourced saliency [4, 10], and reduce training time by supplying additional prior knowledge into the training process [4, 44]. However, a common and valid criticism is the high cost of acquiring human annotations or eye tracking-sourced saliency data. For anomaly detection tasks, such as biometric presentation attack detection (PAD), new attack types (i.e. paper print outs, doll eyes, textured contact lens) are developed frequently. Medical diagnostics may require continual update of human perceptual understanding of anomalies present in medical samples. In all these tasks, requiring few-shot learning solutions functioning well in an open-set classification regime, collecting more training samples and/or more human saliency data is difficult or even impossible. This is why an effective use of the existing, although always-limited human saliency information originating from domain experts is crucial. An associated question is how and where to appropriately incorporate human perceptual understanding of a task into model training to deliver strong priors allowing for better generalization of such models.
@@
This paper introduces MENTOR, a framework for human perception-guided pretraining, evaluated in the context of open-set iris PAD, synthetic face detection, and disease classification from chest X-ray scans. MENTOR relies on the intuition that (as children) we first learn the representations of the visual world without knowing the “class labels,” and we assign meaning in what we see later in our lives. Around being one year old, infants began to point, imitate, and understand objects merely by observation or visual salience, without any explicit knowledge of the object’s application or type (classification) [16]. We emulate this iterative process by first training an autoencoder to learn the associations between an input image and a human saliency map (only visual salience, with no class labels). Once this representation is learned, the decoder-side is removed, a classifier is put on top of the encoder’s latent space, and the entire model is then fine-tuned in a regular way using cross-entropy loss to solve a classification task at hand. Fig. 1 illustrates this pipeline. Our experiments show that such class label-agnostic, but human perception-aware pretraining of the model’s backbone allows for (a) better generalization to unknown samples than models fine-tuned in the same way but initialized with weights obtained in pretraining on a massive amount data such as ImageNet, (b) same as (a) when compared to models trained with a state-of-the-art human perception guidance incorporated into a loss function. We make these observations for three different domains, in which humans can deliver meaningful saliency information (iris PAD, synthetic face detection, and chest X-ray-based diagnosis), and for three different neural network architectures (ResNet, Inception and EfficientNet). In other words, MENTOR efficiently incorporates the limited human perceptual understanding of a domain into training of convolutional neural networks (CNN).
@@
Two core differences between MENTOR and state-of-the-art unsupervised representation learning approaches, such as DINO [5], are that (a) MENTOR leverages human perception-based guidance in the pre-training phase, and (b) MENTOR operates in a very data-limited regime, since it’s targeted to domains, in which collecting data is intricate, but humans competent in a given domain can be found (i.e. physicians in medical diagnosis). It is also noteworthy that MENTOR works with any CNN architecture, and without any new or invasive architectural changes. This paper is organized around the following two research questions:
@@
  • RQ1: Does MENTOR improve the performance of iris presentation attack detection, synthetic face detection, and chest X-ray-based diagnosis compared to conventional vision pretraining techniques [13] and human perception-guided methods? (answered in Sec. 5.1)
  • RQ2: Can MENTOR be applied to existing human perception-guided techniques to further increase their generalization? (answered in Sec. 5.2)
The sources codes of the proposed approach are offered at https://github.com/CVRL/MENTOR
@@ 2.1 Human Saliency-Guided Training
The most architecture-agnostic implementations of incorporating human saliency into model’s training are those using perception-specific loss function components. Boyd et al. [4] introduced the CYBORG loss function, which penalizes the model for divergence between human saliency maps and model’s Class Activation Maps (CAM) [48]. While this method aligns the activations of the model’s last convolutional layer with human saliency, it’s unclear if that mechanism actually builds a human perceptual intelligence into the model, especially in its earlier layers, or acts as a regularizer by lowering the entropy of model’s salience [37]. Other methods use autoencoding strategies, such as UNET+Gaze to incorporate human salience into training a classifier, though these operate as somewhat independent tasks [22]. These methods train the classification and human-salient loss components jointly, which may often conflict with one another during training, leaving to information collapse during the reconstruction phase, or forcing the model to obey only one loss term at a time. Instead, the MENTOR approach proposed in this paper is different from UNET+Gaze and CYBORG (as well as other similar human perception-based training strategies) in a sense that it encourages the model to first build associations between input images and human-sourced salient features, without telling the model what task these features are useful for, to create more general interpretations of the visual world. This allows the two tasks to be fully disentangled so that generalization capabilities can be built upon existing feature representations learned from the previous training step.
@@ 2.2 Efficient Use of Human Annotations
Crum et al. [8] applied human salience in a Teacher-Student training paradigm to make a better use of limited human salience data. In their work, human annotations were first used to train “teacher” models, which are then used to generate subsequent model salience for training “student” models. Teachers were trained using the CYBORG loss, and afterwards used to generate saliency maps on un-annotated training samples using CAM [48] and RISE [35]. This training paradigm boosted performance in synthetic face detection and iris PAD. Interestingly, the MENTOR’s byproduct is an autoencoder that generates human saliency maps for unseen data, and thus it can complement approaches such as the above teacher-student learning paradigms. However, challenges with the Teacher-Student paradigm involve which is the optimal “teacher” model to select, and obtaining sufficient human-salient training samples for the Teachers to learn the task and subsequently train Student models.
@@ 3.1 Methodology
MENTOR is a novel approach of incorporating human salience into model training through a two-part training series. First, an autoencoder is trained to generate human-like saliency given an input image, but without any class labels (“Step 1” in Fig. 1). In that way, the model creates associations between input samples and human-salient regions. Once these associations are made, we put a single, fully-connected layer (initialized with random weights) on top of the encoder’s embeddings and fine-tune the entire architecture to solve the classification task (see “Step 2” in Fig. 1).
It’s analogous to theories involving how humans gain perceptual understanding of the visual world, by first being exposed to stimuli without being given explanations (or “class labels”) of what they see. As shown in Sec. 5), such human perception-based pre-training results in a better performance in three anomaly detection tasks (from different domains) compared to those observed for models initialized with weights obtained after training with a large visual dataset (ImageNet), and than for models trained with state-of-the-art loss function-based human perception guidance.
@@
where $y_{k}$ is a class label for the $k$-th sample, $\bm{1}$ is an indicator function equal to $1$ when $y_{k}\in\mathcal{C}_{c}$, and 0 otherwise, $C$ is the total number of classes, $K$ is the number of samples in a batch, $\textbf{s}_{k,\mathcal{C}_{c}}^{\text{human}}$ and $\textbf{s}_{k,\mathcal{C}_{c}}^{\mathcal{F}}$ are the human and the model $\mathcal{F}$’s saliency for the $k$-th sample representing class $\mathcal{C}_{c}$, respectively, and finally $\mathcal{D}$ is the measure of dissimilarity between saliency maps (e.g. the Mean Square Error distance).
@@ 3.2 Neural Network Architectural Choices
The MENTOR approach does not specify a type of autoencoder. We experimented with UNET [39] and UNET++ [49] with three different backbones: ResNet152 [17], Inception-V4 [40] and EfficientNet-b7 [41]) using the same training configurations [19] to observe stability of the proposed approach across visual domains and architectures.
Specific domains and backbones for the autoencoder reported in Sec. 5 are as follows: UNET (all backbones for iris PAD domain) and UNET++ (all backbones for synthetic face detection and chest X-ray-based diagnosis).
Also, MENTOR does not specify the classifier put in Step 2 on top of the encoder embeddings. We demonstrate effectiveness of this approach by adding a single-layer classifier to minimize the extra architectural components added to human saliency-guided pre-trained encoder, but there are no theoretical reasons for not exploring the generalization gain achieved for deeper structures as a future research.
@@ 3.3 Training and Evaluation Details
MENTOR does not need any specific hyperparameter settings. For reference purposes, we provide settings we have used in this work. The autoencoders in Step 1 were trained with the AdamW optimizer, with a learning rate of 0.0001, and with a batch size of 8. MSE loss was used to assess the agreement bewteen human saliency and the predicted saliency maps, all scaled to a canonical $224\times 224$ pixel resolution. Training was continued until 50 epochs, but many models converged quickly (in less than 5 epochs). All full models in Step 2 were trained using cross-entropy loss using Stochastic Gradient Descent (SGD), with a learning rate of 0.005 decreased by 0.1 every 12 epochs. As in Step 1, maximum number of epochs was 50. CYBORG and UNET+Gaze models were trained from the source codes and training configurations provided by their work [4, 22].
We evaluate the proposed MENTOR approach using Area Under the Receiver Operating Characteristic Curve (AUROC). For all experiments we perform ten independent training runs instantiated using different seeds.
@@ 4.1 Training and Validation Data Subsets
For training the subsequent iris PAD classification models (“Step 2” in Fig. 1), the same 765 images (but with class labels and no salience maps) were used as training images, with additional 23,312 validation images sampled from the same superset as used in [2]. The train and validation sets are subject-disjoint. The 765 training samples were comprised of 198 live irises and 567 spoof irises.
@@
We trained the autoencoder using real and synthetically-generated face images and human saliency maps offered by [4]. 363 humans recruited via Amazon Mechanical Turk annotated regions of the image deemed important to their classification decision (real / fake). Only correctly classified samples (images and accompanying saliency maps) were used during training. For simplicity, 765 samples (379 real / 386 synthetic) were randomly selected without replacement to match the number of samples available for Domain 1 (Iris PAD).
@@
The autoencoder was trained with normal and abnormal chest X-ray images [21], and corresponding eye-tracking visual salience collected from radiologists examining these X-ray scans [1]. We consider only abnormal categories represented by at least 200 training images, that is: atelectasis, cardiomegaly, edema, lung opacity, pleural effusion, pneumonia, and support devices. We removed images containing erroneous class labels and consider only images where radiologists were certain of their classification decision. The intensity of pixels in eye tracking-based salience maps is proportional to the strength of the eye fixation. If a sample included human saliency maps from multiple radiologists, we combined these maps by taking the maximum intensity for each pixel. This was done to preserve salient information provided by each expert, who may consider different features when classifying the samples, instead of considering only the most “popular” features. All images were resized to $224\times 224$. For consistency, 765 samples (223 normal / 542 abnormal) and accompanying saliency maps were used during the training phases to match with sample sizes used in two other domains.
For training the chest X-ray classification models (“Step 2” in Fig. 1), all 765 images were used as training images, in conjunction with additional 4,878 subject-disjoint validation samples sampled from the same dataset.
@@ 5 Results
Table 1: Mean and standard deviations of AUROC scores (calculated independently for each combination of domain-architecture) across 10 independent runs. The indicates that no human saliency was used during training, whereas indicates human saliency was used during training.
Dataset Cross-entropy CYBORG [4] UNET+Gaze [22] MENTOR (ours)
       Trained Using Human Saliency
@@
         ResNet152 0.879$\pm$0.02 0.897$\pm$0.02 0.892$\pm$0.03 0.912$\pm$0.02
         Inception-V4 0.879$\pm$0.03 0.794$\pm$0.04 0.807$\pm$0.024 0.899$\pm$0.05
         EfficientNet-7b 0.881$\pm$0.01 0.801$\pm$0.03 0.830$\pm$0.039 0.901$\pm$0.02
@@
         ResNet152 0.540$\pm$0.04 0.508$\pm$0.05 0.432$\pm$0.042 0.538$\pm$0.05
         Inception-V4 0.519$\pm$0.06 0.484$\pm$0.03 0.438$\pm$0.053 0.585$\pm$0.03
         EfficientNet-7b 0.465$\pm$0.02 0.500$\pm$0.03 0.453$\pm$0.086 0.565$\pm$0.03